OpenStudy (anonymous):

I need to make this all one log function: 3logx−2log(x^2+1)+2log(x−1) i got: log((x^3)/(x^2+1)^2*(x-1)^2) but that's wrong. ><

OpenStudy (anonymous):

\[ 3\log x−2\log (x^2+1)+2 \log(x−1) = \log \frac{x^3(x-1)^2}{(x^2+1)^2}\]

OpenStudy (anonymous):

i'm pretty sure I tried that as well as one of my options. we're allowed to answer this question as many times as we want. It didn't work. ><

OpenStudy (anonymous):

Do you know the answer?

OpenStudy (anonymous):

No, we aren't told it. It just says whether we're right or not.

OpenStudy (anonymous):

\[ 3\log x - 2\log(x^2+1)+2\log(x-1)\\ \log x^3-\log(x^2+1)^2+\log(x-1)^2\\ \log\frac{x^3(x-1)^2}{(x^2+1)^2} \]Same as FFM's solution.

OpenStudy (anonymous):

the problem then could be written wrong. i don't see anything wrong with these solutions posted...

OpenStudy (anonymous):

yeah i dunno. i checked my working and stuff and did it again. but the system still called it wrong. it's okay. thanks though.